# JammingSnake: A follow-the-leader continuum robot with variable stiffness based on fiber jamming

### v1

@@ V Discussion
    Reasonable Actuation Time of Stiffness Variation: Our prototype achieves a fast actuation time of 2 seconds, which is significantly competitive. Some other designs, like those using SMP [30] and LMPA [15], have actuation times that can be much slower. Our faster response time can be critical for applications that demand real-time adjustments where delays could impact efficiency and safety. In addition, the current actuation time is limited by the actuator, not the mechanism, making it easier to be reduced.
@@
    Error Resilience Over Long Distances: Our prototype employs a decoupled control mechanism, avoiding the length-dependent errors seen in other designs. Coupled control systems, which link robot movement and stiffness variation through shared actuation like tendon-pulling, often cause unintended interactions—such as unintentional bending during stiffness adjustments or disrupted stiffness during bending. These interactions complicate precise control, increasing operational errors, especially in tasks requiring the robot as an end effector. For example, several designs (e.g., [29, 5, 11]) exhibit length-dependent errors, limiting their effectiveness as end-effectors over longer distances. Our prototype, with its decoupled mechanism, ensures consistent performance and accuracy regardless of the robot’s length.

### v2

@@ V Discussion
    Reasonable Actuation Time of Stiffness Variation: Our prototype achieves a fast actuation time of 2 seconds, which is significantly competitive. Some other designs, like those using SMP [30] and LMPA [15], have actuation times that can be much slower. Our faster response time can be critical for applications that demand real-time adjustments where delays could impact efficiency and safety. In addition, the current actuation time is limited by the actuator, not the mechanism, making it easier to be reduced to less than 0.5 s.
@@
    High Shape Conversation Performance: Our prototype employs a decoupled control mechanism, avoiding the posture errors seen in other designs during the stiffness variation procedure. Coupled control systems, which link robot movement and stiffness variation through shared actuation like tendon-pulling, often cause unintended interactions such as unintentional vibration of the robot during stiffness adjustments or disrupted stiffness during bending. For instance, HARP [11] frequently exhibits end-effector shaking during propagation, worsening as its length increases. Such interactions complicate precise control, increasing posture errors. In contrast, our decoupled mechanism minimizes shape changes during propagation.